Each side of a marble tile is 16 cm long.How many tiles are needed to cover an area of 5120 square cm?
Found 2 solutions by Cromlix, killmole1:
Answer by Cromlix(4381)

Area of tile = 16cm * 16cm = 256cm^2
5120cm^2 / 256cm^2 = 20
You will require 20 tiles to
cover 5120cm^2.
Hope this helps.
:-)

Answer by killmole1(1)

A=l*W
A=5120cm^2/16cm=16cm/16cm W
=320 m width